sir n cube greater than
=n. what is the relation

Grade:11

1 Answers

Mohan Chandar
9 years ago
We shall prove it by math Induction... When n=1, n^3=1^3=1=n... Let K^3 >= k... we shall prove that (K+1)^3 >= k+1... (k+1)^3 = K^3 + 1 + 3k^2 + 3k >= K + 3(k+1) + 1 >= 3(k+1) + k+1 >= 4(k+1) >= k+1....... Hence proved....